OpenStudy (anonymous):

Can I get some help simplifying a few problems? Such as 1) sqrt 108x^3y^2

OpenStudy (agreene):

So, with this you can only pull out the constant, because the variables can be values that would make splitting them erroneous. \[\sqrt{108x^3y^2}=\sqrt{108}\sqrt{x^3y^2}\] try simplifying that

OpenStudy (anonymous):

Alright, one moment.

OpenStudy (agreene):

btw, if we know that x,y are positive, we could split it a bit more. if they do become negative the imaginary parts are what keep us from splitting them

OpenStudy (anonymous):

I do know that with a square root, by the way I was taught would lead to x^3 to being something like this: \[xy \sqrt{108x}\]

OpenStudy (agreene):

right, it is correct to lead polys with the highest exponents first. in this case, we just pull the constant out and simplify that, since that's all we can really do

OpenStudy (anonymous):

So, would that be this? : \[6xy \sqrt{3x}\]

OpenStudy (agreene):

right, \[xy(108x)^{\frac12}=6xy(3x^{\frac12})\]

OpenStudy (anonymous):

OH, got it.

OpenStudy (agreene):

oh, sorry, i just default to doing that--i havent written a sqrt sign in years.. \[\sqrt{x}=x^{\frac12}\]

OpenStudy (anonymous):

1/2 is just simply saying 'The square root.'

OpenStudy (anonymous):

Um, would you mind helping me with my other problems?

OpenStudy (agreene):

right, and cubed root is 1/3 fourth root is 1/4

OpenStudy (agreene):

sure.

OpenStudy (anonymous):

The next one up is \[\sqrt{-162}\]

OpenStudy (agreene):

so the negative just tells us there will be an i in front. so whats \(\sqrt{162}\) ? then just put an i in front

OpenStudy (anonymous):

So, \[9\sqrt{2i}\]?